MAT 142, HONORS CALCULUS, SUNY AT STONY BROOK, FALL 2001

Lecturer: Prof. Christopher Bishop, bishop@math.sunysb.edu, office is 4-112 in Math Building. Office hours are M TH 1:15-2:15.

Text: Calculus, 10th edition, Thomas and Finney.

Syllabus: We will cover approximately Chapters 5,6,7 and 8 with certain omissions and additions of sections. A more detailed syllabus will be provided.

Classdays. Classes begin on Monday, August 27 and end on Thursday, December 13.

Holidays: There are no classes on Labor Day (Monday, September 3), Rosh Hashana (Tuesday and Wednesday, September 18-19), Yom Kipur (Thursday, September 27) and Thanksgiving (November 19-25).

Lecture times: Tuesday and Thursday, 2:20-3:40 pm.

Lecture location: Room 231 in the Stony Brook Union.

Recitation times and locations: There are two recitiations.

• Section 1 meets Monday at 12:40-2:00 in ESS 177. The TA is Chanyoung Sung. His office number is ??? in the Math Building and his email is cysung@math.sunysb.edu. His office hours are ?????????????????????.

• Section 2 meets Friday at 12:40-2:00 in ESS 177. Chanyoung Sung is also the TA for this section.

Homework. A list of problems from the text and supplementary material will be handed out. The recitation instructor will assign, collect and return homework. He may also give weekly quizzes based on the assigned problems.

Examinations. There will be two midterm tests: Thursday, October 11 and Thursday, November 15. These will be given during the regular lecture time. The final exam is on Tuesday, December 18 at 2:00-4:30pm. The location will be announced in class during the last week of the semester.

Grades. Grades are computed using: Recitation 10%, Midterms 25% each, Final 40%.

Disabilities. If you have any condition, such as physical or mental disability, that will make it difficult to complete the course or that will require extra time on examinations, notify your lecturer during the first two weeks of class so that appropriate arrangements can be made.

Math Learning Center: (Physics A-125/127) This is located on the A level of the physics building and provides help with pre-calculus material. The same room also contains the Calculus Resource Room where you can go for help with Calculus classes. It is staffed most days and some evenings; an exact schedule will be posted in the room. Note that this room is closed during final exam week. It also gets crowded before exams, so don’t wait until the last minute to get help there.

Calculus Online: There are many math sites on the internet. For information on the history of Calculus and short biographies of famous mathe-

maticians, try http://www-groups.dcs.st-and.ac.uk/~history and click

on History Topics or the Biography Index. You can find information about math classes at Stony Brook at www.math.sunysb.edu

Drop dates: Last day for Add/Drop is Monday, September 10. Last day to switch to lower numbered math course is Thursday, October 11. The last day to withdraw from a course is Friday, November 2.